OpenStudy (anonymous):

tell me how you get the following thing for this shape - apothem -radius -1/2 Central angle -area

OpenStudy (anonymous):

please give me some number answers too,..

OpenStudy (anonymous):

we're basically solving for the parts of that triangle...

OpenStudy (anonymous):

the radius is the hypotenuse of the triangle the apothem is the distance from the center to the midpoint of the side. (in the triangle, it's the intermediate length side.)

OpenStudy (campbell_st):

an apothem is a line segment from the midpoint if a side of a polygon to the centre

OpenStudy (anonymous):

oh yeah.... line segment... what he ^^^ said... :)

OpenStudy (anonymous):

so uum.. how do i find what im asking...

OpenStudy (anonymous):

each side is 8 so half of it is...??? that will be your shortest side of the triangle.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if you draw a segment from each vertex of the polygon to the center, how many triangles you'd get?

OpenStudy (campbell_st):

the half angle the polygon is regular so the angle at the centre 360/6 = 60 half angle = 60/2 = 30

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

how do i find the area?

OpenStudy (anonymous):

so now you know what kind of triangle we're dealing with.. a 30-60-90 triangle

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

we're getting to the area... all we need is the intermediate side length.. because area of the triangle is (1/2)b*h, right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

 because this is 30-60-90, do you know what the ? without using the pythagorean theorem?